OpenStudy (anonymous):

A car is traveling on a highway. The distance (in miles) from its destination and the time (in hours) is given by the equation d=420-65t. What is the d-intercept of the line? (Remember the slope-intercept form is y=mx+b) a. (0, 420) b. (420, 0) c. (0, 65) d. (65, 0)

OpenStudy (anonymous):

I'll Try My Best, Hold On.

OpenStudy (anonymous):

Got (C). But I may Be Wrong.. No Math Wizard.

OpenStudy (phi):

they want you to "match up" d=420-65t with y=mx+b if you re-arrange the first equation to this d = -65 t + 420 and match it to y = m x + b you see d matches up with "y" d-intercept would be translated as y-intercept b is the y intercept. what matches with b ?

OpenStudy (anonymous):

Thanks a million!